1. In this problem, x = c1 cos t + c2 sin t is a two-parameter family of solutions of the second-order DE x'' + x = 0. Find a solution of the second-order IVP consisting of this differential equation and the given initial conditions.

x(pi/3) = square root of 3 / 2, x prime (pi/3) = 0

x = ____
1
Expert's answer
2020-06-16T15:54:40-0400

 x'' + x = 0

"\frac {d^2x}{dt^2} + x = 0"

Let x = emt be a solution

So m²emt+ emt = 0

=> emt(m²+1) = 0

Auxiliary equation is m²+1=0 as emt ≠0

=> m = ±i

So the general solution of the given differential equation is

x = (C1 cost + C2 sin t)e0*t

=> x = C1 cost + C2 sin t

Differentiating with respect to t

"\frac {dx}{dt}" = - C1 sin t + C2 cos t

Using the initial conditions

x = "\frac {√3}{2}" when t = "\frac {π}{3}"

"\frac {dx}{dt}" = 0 when t = "\frac {π}{3}"

So C1 cos("\frac {π}{3}" ) + C2 sin("\frac {π}{3}" ) = √3/2

=> "\frac {1}{2}" C1 + "\frac {√3}{2}" C2 = "\frac {√3}{2}"

=> C1+ √3C2 = √3 -----------eq(1)

Again "\frac {dx}{dt}" = 0 when t = "\frac {π}{3}"

So,. - C1 sin("\frac {π}{3}" ) + C2 cos("\frac {π}{3}" ) = 0

=> - "\frac {√3}{2}" C1 + "\frac {1}{2}" C2 = 0

=> -√3C1 + C2 = 0

Multiplying by √3

-3C1 + √3C2 = 0 --------------eq(2)

Subtracting eq(2) from eq(1) we get

4C1 = √3

So C1 = "\frac {√3}{4}" and hence C2 = √3C1= "\frac {3}{4}"

So the actual solution is

x = "\frac {√3}{4}" cos t + "\frac {3}{4}" sin t